Consider the following unbalanced equation for the combustion of hexane:
αC6H14(g)+βO2(g)→γCO2(g)+δH2O(g)

A. Balance the equation.

B. Determine how many moles of O2 are required to react completely with 5.2 moles C6H14
Answer #1

A. To balance the equation:

- Start by balancing the elements that exist in one reactant and one product. You can start by carbon or hydrogen but not oxygen (O exist in one reactant O2 but in 2 products (CO2 and H2O))

C6H14 + O2 \rightarrow6 CO2 + 7 H2O

- Balance the other elements:

In the right: 6x2 + 7x1 = 19 O

\Rightarrow U need 19 O in the left : (19/2) O2

C6H14 + (19/2)O2 \rightarrow 6 CO2 + 7 H2O

- Delete the fractions by multiplying both sides by the smallest integer number to obtain whole numbers for all reactatnts and products:

2x(C6H14 + (19/2)O2 \rightarrow 6 CO2 + 7 H2O)

\Rightarrow2 C6H14 + 19 O2\rightarrow12 CO2 + 14 H2O

- Check the balance of elements on the right and left:

LEFT RIGHT

2x6 C = 12 C 12 C

2x14 H = 28 H 14x2 H= 28 H

19x2O= 38 O 12x2 + 14 = 38 O

The balanced equation is:

2 C6H14 + 19 O2\rightarrow12 CO2 + 14 H2O

B. Based on the stoichiometry of the reaction:

\frac{n_{C_6H_{14}}}{2} =\frac{n_{O_2}}{19}

\Rightarrown_{O_2}= \frac{19\times n_{C_6H_{14}}}{2}= \frac{19\times 5.2}{2}= 49.4 mol